Grade 12th PassMagical Mathematics[Interesting Approach]

If 8 sin2 A + 4 cos2 A = 5, THEN 4 sin2 A + 8 cos2 A is equal to?

Vikas TU
15 Years agoGrade 12th Pass

1 Answer

vikas askiitian expert
15 Years ago

sin2A = 1-cos2A         ............1

8sin2A + 4cos2A = 5          ..........2       (given)

putting 1 in 2

8 - 8cos2 + 4cos2A = 5

cos2A = 3/4

sin2A = 1-cos2A =  1/4

so , value of 4sin2A+8cos2A = 4(1/4) + 8(3/4)

             =7